the inverse demand functions of two towns for a product are given as
( Q1/2)^1/4=P and (2/10-Q1/10)=P find the market demand of these two towns
1
Expert's answer
2020-03-25T10:43:35-0400

Market demand is the summation of the individual demand curves. (Q1/2)1/4= P

Re-arranging in terms of Q, we get by taking exponential power on the other side,

Q1/2 = P4

Q1 = 2P4

(2/10-Q1/10)=P "\to" 10P = 2-Q1 → Q1 = 2 - 10P

2P4 = 2 - 10P

P4 = 1 - 5P

P4 + 5P = 1